OpenStudy (anonymous):

A record collection consists of eight rock albums, twelve blues albums, and four classical albums. If two records are chosen random, find the probability that they are both blues?

OpenStudy (anonymous):

\[\frac{12}{24}\times \frac{11}{23}\]

OpenStudy (anonymous):

=\frac{11}{46}\]